5x+4y=19.75
x+2y=5.75

Answer by mananth(16946)

5.00 x + 4.00 y = 19.75 .............1

1.00 x + 2.00 y = 5.75 .............2
Eliminate y
multiply (1)by 1.00
Multiply (2) by -2.00
5.00 x 4.00 y = 19.75
-2.00 x -4.00 y = -11.50
Add the two equations
3.00 x = 8.25
/ 3.00
x = 2.75
plug value of x in (1)
5.00 x + 4.00 y = 19.75
13.75 + 4.00 y = 19.75
4.00 y = 19.75 -13.75
4.00 y = 6.00
y = 1.50
x=3,y=1.5